# Pointwise convergence of ergodic averages with Möbius weight

### v1

@@ 5.2. A lacunary subsequence trick
We use a lacunary subsequence trick in the proofs of our pointwise convergence results. Such a trick roughly states that if $A_{N}\colon X\to\mathbb{C}$ are some measurable functions and we have strong quantitative decay for $\|A_{N_{j}}\|_{L^{1}(X)}$ for some lacunary sequence $(N_{j})_{j\in\mathbb{N}}$, then provided that $A_{N}$ does not vary too much on intervals of the form $[N_{j},N_{j+1}]$ in $L^{1}(X)$ norm, the sequence $A_{N}$ must converge to 00 in $L^{\infty}(X)$ norm. Variants of this idea are frequently used to establish convergence of ergodic averages; see for example [10, Section 5].
@@
Let $B>1$ and $c>0$. Let $(X,\nu)$ be a probability space, and for $N\geq 1$ let $A_{N}\colon X\to\mathbb{C}$ be a measurable function.
  1. (i)
    If for any $N\geq 3$ we have
@@
    and
    (5.2) $\left\|\sup_{M\in[N,(1+(\log N)^{-(B/2-1)})N]}|A_{M}-A_{N}|\right\|_{L^{1}(X)}\ll(\log N)^{-B/3},$
    then, for any $\varepsilon>0$, we have $\lim_{N\to\infty}(\log N)^{B/4}\|A_{N}\|_{L^{\infty}(X)}=0$.
  2. (ii)
    If for any $N\geq 3$ we have
    (5.3) $\|A_{N}\|_{L^{1}(X)}\ll(\log\log N)^{-c}$
    and
    (5.4) $\begin{split}\left\|\sup_{M\in[N,\exp((\log N)^{1+(\log\log N)^{-c}})]}|A_{M}-A_{N}|\right\|_{L^{1}(X)}\ll(\log\log N)^{-c},\end{split}$
    then we have $\lim_{N\to\infty}(\log\log N)^{c/3}\|A_{N}\|_{L^{\infty}(X)}=0$.
  3. (iii)
    If for any $N\geq 3$ we have
    (5.5) $\|A_{N}\|_{L^{1}(X)}\ll(\log\log N)^{-c},$
    then, for any sequence $(N_{j})_{j\in\mathbb{N}}$ satisfying $N_{j}\gg\exp(\exp(j^{2/c}))$, we have
    $\lim_{j\to\infty}(\log\log N_{j})^{c/3}\|A_{N_{j}}\|_{L^{\infty}(X)}=0$.
@@
(i) Let $\eta>0$ be a small enough constant, and set $N_{j}=\exp(\eta\cdot j^{2/B})$ for all $j\in\mathbb{N}$. Note that $N_{j+1}\leq N_{j}(1+(\log N_{j})^{-(B/2-1)})$ for all large enough $j\in\mathbb{N}$. For $\delta>0$, denote
(5.6) $\begin{split}E(\delta)&=\{x\in X\colon\limsup_{N\to\infty}\,(\log N)^{B/4}|A_{N}(x)|\geq\delta\},\\ E_{j}(\delta)&=\{x\in X\colon|A_{N_{j}}(x)|\geq\frac{\delta}{3}(\log N_{j})^{B/4}\},\\ E_{j}^{\prime}(\delta)&=\{x\in X\colon\sup_{M\in[N_{j},N_{j+1}]}|A_{M}(x)-A_{N}(x)|\geq\frac{\delta}{3}(\log N_{j})^{B/4}\}.\end{split}$
Then we have
(5.7) $E(\delta)\subset\bigcap_{i=3}^{\infty}\bigcup_{j\geq i}(E_{j}(\delta)\cup E_{j}^{\prime}(\delta)).$
By Markov’s inequality, (5.1) and (5.2), we have the bounds
$\nu(E_{j}(\delta))$ $\ll_{\delta}(\log N_{j})^{B/4-B}\ll j^{-3/2},$
$\nu(E_{j}^{\prime}(\delta))$ $\ll_{\delta}(\log N_{j})^{B/4-B}\ll j^{-3/2}.$
Since $\sum_{j\geq 3}j^{-3/2}<\infty$, from (5.7) and the Borel–Cantelli lemma we conclude that $\nu(E(\delta))=0$. By the countable additivity of $\nu$, this then implies that
$\nu(\bigcap_{\delta>0}E(\delta))=\nu(\bigcap_{j\geq 1}E(1/j))=0,$
as desired.
(ii) Let $\eta>0$ be a small enough constant, and set $N_{j}=\exp(\exp(\eta\cdot j^{2/c}))$ for all $j\in\mathbb{N}$. Then we have $N_{j+1}\leq\exp((\log N_{j})^{1+(\log\log N_{j})^{-c/2}})$ for all large enough $j\in\mathbb{N}$. For $\delta>0$, denote
$E(\delta)$ =$=\{x\in X\colon\limsup_{N\to\infty}\,(\log\log N)^{c/3}|A_{N}(x)|\geq\delta\}$
$E_{j}(\delta)$ =$=\{x\in X\colon|A_{N_{j}}(x)|\geq\frac{\delta}{3}(\log\log N_{j})^{-c/3}\},$
$E_{j}^{\prime}(\delta)$ =$=\{x\in X\colon\sup_{M\in[N_{j},N_{j+1}]}|A_{M}(x)-A_{N}(x)|\geq\frac{\delta}{3}(\log\log N_{j})^{-c/3}\}.$
Then we have (5.7). From (5.3), (5.4) we have
$\nu(E_{j}(\delta)),\nu(E_{j}^{\prime}(\delta))\ll_{\delta}(\log\log N_{j})^{-2c/3}\ll j^{-4/3}$
for all $j\geq 3$, so as above the Borel–Cantelli lemma gives $\nu(\bigcap_{\delta>0}E(\delta))$, which was the claim.
(iii) Let a sequence $N_{j}\gg\exp(\exp(j^{2/c}))$ be given. For $\delta>0$, denote
$E(\delta)$ =$=\{x\in X\colon\limsup_{i\to\infty}\,(\log\log N_{i})^{c/3}|A_{N_{i}}(x)|\geq\delta\},$
$E_{j}(\delta)$ =$=\{x\in X\colon|A_{N_{j}}(x)|\geq\frac{\delta}{2}(\log\log N_{j})^{-c/3}\}.$
Then we have
$E(\delta)\subset\bigcap_{i=1}^{\infty}\bigcup_{j\geq i}E_{j}(\delta).$
By Markov’s inequality and (5.5), we have
$\nu(E_{j}(\delta))\ll_{\delta}(\log\log N_{j})^{-2c/3}\ll j^{-4/3}.$
Since $\sum_{j\geq 2}j^{-4/3}<\infty$, the Borel–Cantelli lemma again gives us $\nu(\bigcap_{\delta>0}E(\delta))=0$, as desired. ∎

### v2

@@ 5.1. The Möbius function
5.2. Multiplicative functions satisfying the Siegel–Walfisz condition
Our goal in this subsection is to show that any multiplicative function satisfying the Siegel–Walfisz assumption (Definition 1.5) is close to a function whose $u^{k}[N]$ norm decays faster than any power of logarithm.
Proposition 5.4.
Let $k\in\mathbb{N}$ and $A>0$. Let $g\colon\mathbb{N}\to\mathbb{C}$ be a multiplicative function satisfying the Siegel–Walfisz property. Then we have a decomposition $g=g_{1}+g_{2}$ with $g_{1},g_{2}\colon\mathbb{N}\to\mathbb{C}$ satisfying $|g_{1}|,|g_{2}|\leq|g|$ and such that for $N\geq 3$ we have
$\|g_{1}\|_{u^{k}[N]}\ll_{A,k}(\log N)^{-A}.$
and for $r\geq 1$ we have
$\mathbb{E}_{n\leq N}|g_{2}(n)|^{r}\ll_{r}(\log N)^{-1+o(1)}.$
Throughout this section, let
$Q_{n}\coloneqq\exp((\log\log n)^{2}),\quad R_{n}\coloneqq\exp((\log n)/(\log\log n)^{2})$
for $n\geq 3$ and $Q_{1}=Q_{2}=R_{1}=R_{2}=1$, and define the function
(5.1) $\widetilde{g}(n)\coloneqq g(n)(1-1_{p\mid n\implies p\not\in(Q_{n},R_{n})}).$
In other words, $\widetilde{g}$ is the restriction of $g$ to those integers having at least one prime factor from $(Q_{n},R_{n})$. The following lemma shows that the function $\widetilde{g}$ is close to $g$ in $L^{r}[N]$ norm.
Lemma 5.5.
Let $C>0$, $r\geq 1$, and let $g\colon\mathbb{N}\to\mathbb{C}$ satisfy $|g(n)|\leq d(n)^{C}$ for all $n\in\mathbb{N}$. Then for $N\geq 3$ we have
(5.2) $\mathbb{E}_{n\leq N}|g(n)-\widetilde{g}(n)|^{r}\ll_{C,r}(\log N)^{-1+o(1)}.$
Proof.
The left-hand side of (5.2) is by the triangle inequality and the upper bound on $g$ bounded by
$\ll\mathbb{E}_{n\leq N}d(n)^{Cr}1_{p\mid n\implies p\not\in(Q_{N},R_{\sqrt{N}})}+O(N^{-1/2+o(1)})\coloneqq S+O(N^{-1/2+o(1)}).$
Applying Shiu’s bound [31, Theorem 1] followed by Mertens’s theorem, we see that
$S$ $\ll\prod_{p\in[1,N]\setminus(Q_{N},R_{\sqrt{N}})}\left(1+\frac{d(p)^{Cr}-1}{p}\right)\prod_{p\in(Q_{N},R_{\sqrt{N}})}\left(1-\frac{1}{p}\right)$
$=\prod_{p\leq N}\left(1+\frac{2^{Cr}-1}{p}\right)\prod_{p\in(Q_{N},R_{\sqrt{N}})}\left(1+\frac{2^{Cr}-1}{p}\right)^{-1}\left(1-\frac{1}{p}\right)$
$\ll_{C,r}(\log N)^{2^{Cr}-1}\left(\frac{\log Q_{N}}{\log R_{\sqrt{N}}}\right)^{2^{Cr}}$
$\ll_{C,r}(\log N)^{-1+o(1)},$
as desired. ∎
The next lemma shows that the condition on the prime factors in the definition of $\widetilde{g}$ can be replaced with a sieve weight up to a small error.
Lemma 5.6.
Let $N\geq 3$ and $C>0$. There exist real numbers $\lambda_{r}\in[-1,1]$ such that for any $A>0$ we have
(5.3) $\mathbb{E}_{N-N(\log N)^{-A}<n\leq N}\left|1_{p\mid n\implies p\not\in(Q_{n},R_{n})}-\sum_{\begin{subarray}{c}r\mid n\\ r\leq N^{1/10}\end{subarray}}\lambda_{r}\right|\cdot d(n)^{C}\ll_{A,C}(\log N)^{-A/2}.$
Proof.
By splitting intervals into shorter ones if necessary, we may assume that $A$ is large enough in terms of $C$. Let $\lambda_{r}$ be the upper bound linear sieve coefficients of level $N^{1/10}$ and sifting range $(Q_{N},R_{N-N(\log N)^{-A}})$, as defined in [13, Section 12]. From the definition we have $\lambda_{r}\in\{-1,0,+1\}$. Let $\nu(n)=\sum_{r\mid n}\lambda_{r}$. Then the left-hand side of (5.3) is
$\ll(\log N)^{A/2}\mathbb{E}_{N-N(\log N)^{-A}<n\leq N}|\nu(n)-1_{p\mid n\implies p\not\in(Q_{n},R_{n})}|$
$\quad+\mathbb{E}_{N-N(\log N)^{-A}<n\leq N}d(n)^{C+1}1_{d(n)\geq(\log N)^{A/C}}$
$\coloneqq S_{1}+S_{2}.$
Using Shiu’s bound [31, Theorem 1], we see that
$S_{2}\ll(\log N)^{-2A}\mathbb{E}_{N-N(\log N)^{-A}<n\leq N}d(n)^{2C+1}\ll(\log N)^{-2A+2^{2C+1}-1}\ll(\log N)^{-A}$
by the assumption that $A$ is large in terms of $C$.
Since
$\nu(n)\geq 1_{p\mid n\implies p\not\in(Q_{N},R_{N-N(\log N)^{-A}})}\geq 1_{p\mid n\implies p\not\in(Q_{n},R_{n})},$
for $n\in(N-N(\log N)^{-A},N]$, we have
$\mathbb{E}_{N-N(\log N)^{-A}<n\leq N}|\nu(n)-1_{p\mid n\implies p\not\in(Q_{n},R_{n})}|$
$\quad=\mathbb{E}_{N-N(\log N)^{-A}<n\leq N}(\nu(n)-1_{p\mid n\implies p\not\in(Q_{N},R_{N-N(\log N)^{-A}})})$
$\quad\quad+O\left(\sum_{p\in(Q_{N-N(\log N)^{-A}},Q_{N})\cup(R_{N-N(\log N)^{-A}},R_{N})}\frac{1}{p^{2}}\right).$
By the fundamental lemma of sieve theory ([21, Lemma 6.8]) and Mertens’s theorem, this is
$\ll\exp(-(\log\log N)^{2}/20)+(\log N)^{-A/2},$
which suffices. ∎
We are now ready to show that the Siegel–Walfisz property for $g$ implies the same property for $\widetilde{g}$.
Lemma 5.7.
Let $g\colon\mathbb{N}\to\mathbb{C}$ be a multiplicative function satisfying the Siegel–Walfisz property. Then $\widetilde{g}$ satisfies the Siegel–Walfisz property.
Proof.
Since $\widetilde{g}(n)=g(n)-g(n)1_{p\mid n\implies p\not\in(Q_{n},R_{n})}$, it suffices to show that the function
$n\mapsto g(n)1_{p\mid n\implies p\not\in(Q_{n},R_{n})}$
satisfies the Siegel–Walfisz property. By splitting a long interval into shorter ones, it suffices to show that for any large $A>0$ and any $1\leq a\leq q\leq(\log N)^{A}$ we have
$\left|\mathbb{E}_{N-N(\log N)^{-A}<n\leq N}g(n)1_{p\mid n\implies p\not\in(Q_{n},R_{n})}1_{n\equiv a\ (\mathrm{mod}\ q)}\right|\ll_{A}(\log N)^{-A/2}.$
By Lemma 5.6, it suffices to show that for any $|\lambda_{r}|\leq 1$ we have
$\left|\mathbb{E}_{N-N(\log N)^{-A}<n\leq N}g(n)\sum_{\begin{subarray}{c}r\mid n\\ r\leq N^{1/10}\end{subarray}}\lambda_{r}1_{n\equiv a\ (\mathrm{mod}\ q)}\right|\ll_{A}(\log N)^{-A/2}.$
Exchanging the order of summation and applying the triangle inequality, it suffices to show that
$\sum_{r\leq N^{1/10}}\frac{1}{r}\left|\mathbb{E}_{(N-N(\log N)^{-A})/r<m\leq N/r}g(rm)1_{rm\equiv a\ (\mathrm{mod}\ q)}\right|\ll_{A}(\log N)^{-A/2}.$
Let $\mathcal{A}_{N}$ be the set of $r\in\mathbb{N}$ with $\omega(r)\leq(\log\log N)^{3/2}$. Then by Shiu’s bound we have
$\sum_{\begin{subarray}{c}r\leq N^{1/10}\\ r\not\in\mathcal{A}_{N}\end{subarray}}\frac{1}{r}\left|\mathbb{E}_{(N-N(\log N)^{-A})/r<m\leq N/r}g(rm)1_{rm\equiv a\ (\mathrm{mod}\ q)}\right|$
$\ll(\log N)^{2^{C}-1}\sum_{\begin{subarray}{c}r\leq N^{1/10}\\ r\not\in\mathcal{A}_{N}\end{subarray}}\frac{d(r)^{C}}{r}$
$\leq(\log N)^{2^{C}-1}\sum_{r\leq N^{1/10}}\frac{d(r)^{C+1}}{r\cdot 2^{(\log\log N)^{3/2}}}$
$\ll 2^{-(\log\log N)^{3/2}/2},$
say. In view of this, it suffices to show that for all $r\in\mathcal{A}_{N}\cap[1,N^{1/10}]$ we have
$\left|\mathbb{E}_{(N-N(\log N)^{-A})/r<m\leq N/r}g(rm)1_{rm\equiv a\ (\mathrm{mod}\ q)}\right|\ll_{A}(\log N)^{-3A/2-1}.$
By writing the sum over $((N-N(\log N)^{-A})/r,N/r]$ as a difference of two sums, it suffices to show that for $y\in[N^{9/10}/2,N]$ we have
(5.4) $\left|\mathbb{E}_{m\leq y}g(rm)1_{rm\equiv a\ (\mathrm{mod}\ q)}\right|\ll_{A}(\log N)^{-A-1}.$
We can uniquely factorise $m=\ell m^{\prime}$, where $\ell\mid r^{\infty}$ and $(m^{\prime},r)=1$. By multiplicativity of $g$, we then reduce to showing that
(5.5) $\sum_{\ell\mid r^{\infty}}\frac{1}{\ell}\left|\mathbb{E}_{m^{\prime}\leq y/\ell}g(m^{\prime})1_{r\ell m^{\prime}\equiv a\ (\mathrm{mod}\ q)}1_{(m^{\prime},r)=1}\right|\ll_{A}(\log N)^{-3A/2-1}.$
Let $S_{1}$ denote the part of the left-hand side of (5.5) with $\ell<(\log N)^{4A}$, and let $S_{2}$ denote the part with $\ell\geq(\log N)^{4A}$. By Shiu’s bound, we can crudely estimate
[…]
where for the last line we used the simple inequality
$\sum_{p\mid r}\frac{1}{p^{1/2}}\leq\sum_{p\leq\omega(r)}\frac{1}{p^{1/2}}\ll\omega(r)^{1/2}.$
Since $\omega(r)\leq(\log\log N)^{3/2}$ for $r\in\mathcal{A}_{N}$, we see that $S_{2}\ll_{A}(\log N)^{-3A/2-1}$.
The remaining task is to show that $S_{1}\ll_{A}(\log N)^{-3A/2-1}$. For this, it suffices to show that for any integer $1\leq\ell\leq(\log N)^{4A}$ we have
(5.7) $\left|\mathbb{E}_{m^{\prime}\leq y/\ell}g(m^{\prime})1_{r\ell m^{\prime}\equiv a\ (\mathrm{mod}\ q)}1_{(m^{\prime},r)=1}\right|\ll_{A}(\log N)^{-3A/2-2}.$
From Shiu’s bound and the Siegel–Walfisz assumption on $g$, for any integer $1\leq u\leq N^{1/10}$ and any $r\in\mathcal{A}_{N}\cap[1,N^{1/10}]$ we have
(5.8) $\left|\mathbb{E}_{m^{\prime}\leq y/\ell}g(m^{\prime})1_{r\ell m^{\prime}\equiv a\ (\mathrm{mod}\ q)}1_{u\mid m^{\prime}}\right|\ll_{A,C}\min\left\{\frac{(d(u)d(q))^{C}}{u},(\log N)^{-10A}\right\}.$
Substituting the Möbius inversion formula
$1_{(m^{\prime},r)=1}=\sum_{\begin{subarray}{c}u\mid r\\ u\leq(\log N)^{6A}\end{subarray}}\mu(u)1_{u\mid m^{\prime}}+\sum_{\begin{subarray}{c}u\mid r\\ u>(\log N)^{6A}\end{subarray}}\mu(u)1_{u\mid m^{\prime}}$
and (5.8) into (5.7), and estimating $d(q)\ll q^{o(1)}=(\log N)^{o(1)}$, we reduce to showing that
$\sum_{\begin{subarray}{c}u\mid r\\ u>(\log N)^{6A}\end{subarray}}\frac{d(u)^{C}}{u}(\log N)^{2^{C}-1}\ll_{A,C}(\log N)^{-3A/2-3}.$
Estimating crudely using $d(u)^{C}/u\ll(\log N)^{-5A/2}/u^{1/2}$ for $u\geq(\log N)^{6A}$, and recalling that $A$ is large, it suffices to show that
(5.9) $\sum_{u\mid r}\frac{1}{u^{1/2}}\ll\log N,$
say. Since $r\in\mathcal{A}_{N}$, the left-hand side is
$\prod_{p\mid r}\left(1+\frac{1}{p^{1/2}}+\frac{1}{p^{2\cdot 1/2}}+\cdots\right)\ll\exp(O(\omega(r^{1/2}))\ll\exp(O((\log\log N)^{3/4})),$
which suffices. ∎
For the proof of Proposition 5.4 we also need a bilinear estimate for polynomial phases.
Lemma 5.8.
Let $A,C>0$, $s\in\mathbb{N}$, and let $\alpha_{a}$, $\beta_{a}$ be complex sequences with $|\alpha_{a}|,|\beta_{a}|\leq d(a)^{C}$ for $a\in\mathbb{N}$, and with $\alpha_{a},\beta_{a}$ supported on $a\geq\exp((\log\log N)^{2})$. Let $P(y)=\sum_{0\leq j\leq s}c_{j}y^{s}$ be a polynomial with real coefficients, and suppose that
(5.10) $\left|\sum_{ab\leq N}\alpha_{a}\beta_{b}e(P(ab))\right|\geq N(\log N)^{-A}.$
Then there exists an integer $1\leq\ell\ll_{A,C,s}(\log N)^{O_{A,C,s}(1)}$ such that
$\|\ell c_{j}\|\ll_{A,C,s}\frac{(\log N)^{O_{A,C,s}(1)}}{N^{j}}$
for all integers $1\leq j\leq s$.
Proof.
We may assume that $C>0$ is large and that $A>0$ is large in terms of $C$. Write $\alpha_{a}=\alpha_{a}^{(1)}+\alpha_{a}^{(2)}$ and $\beta_{a}=\beta_{a}^{(1)}+\beta_{a}^{(2)}$, where
$\alpha_{a}^{(1)}=\alpha_{a}1_{|\alpha_{a}|\leq(\log N)^{A/10}},\quad\beta_{a}^{(1)}=\beta_{a}1_{|\beta_{a}|\leq(\log N)^{A/10}}.$
Then, since $A$ is large in terms of $C$, we have
[…]
and similarly with $\beta_{n}^{(2)}$ in place of $\alpha_{n}^{(2)}$. Now, applying the decompositions $\alpha_{a}=\alpha_{a}^{(1)}+\alpha_{a}^{(2)}$, $\beta_{b}=\beta_{b}^{(1)}+\beta_{b}^{(2)}$, (5.11) and Cauchy–Schwarz, the assumption (5.10) yields
$\left|\sum_{ab\leq N}\alpha_{a}^{\prime}\beta_{b}^{\prime}e(P(ab))\right|\gg N(\log N)^{-A/2},$
where $\alpha_{a}^{\prime}=\alpha_{a}^{(1)}/(\log N)^{A/10}$, $\alpha_{b}^{\prime}=\alpha_{b}^{(1)}/(\log N)^{A/10}$. Since $\alpha_{a}^{\prime},\beta_{b}^{\prime}$ are $1$-bounded, the result follows e.g. from [27, Proposition 2.2] (which is an exponential sum estimate over short intervals; weaker results would also suffice). ∎
Proof of Proposition 5.4.
Recall the definition of $\widetilde{g}$ from (5.1). We take $g_{1}=\widetilde{g}$, $g_{2}=g-\widetilde{g}$. Then we immediately have $|g_{1}|,|g_{2}|\leq|g|$. In view of Lemma 5.5, it suffices to show that $\|g_{1}\|_{u^{k}[N]}\ll_{A,k}(\log N)^{-A}$.
Let $C>0$ be such that $|g(n)|\leq d(n)^{C}$ for all $n$. By splitting into short intervals, it suffices to show that for any $A>0$ and any polynomial $P(y)=\sum_{1\leq j\leq k-1}c_{j}y^{j}\in\mathbb{R}[y]$ we have
(5.12) $\left|\mathbb{E}_{N-N(\log N)^{-A}<n\leq N}\widetilde{g}(n)e(P(n))\right|\ll_{A}(\log N)^{-A/2}.$
Let
$\widetilde{g}_{N}(n)=g(n)(1-1_{p\mid n\implies p\in(Q_{N},R_{N})}).$
Write $\mathcal{J}_{N}=(Q_{N-N(\log N)^{-A}},Q_{N}]\cup(R_{N-N(\log N)^{-A}},R_{N}]$ for brevity. Then for $n\in(N-N(\log N)^{-A},N]$ we have
$\widetilde{g}(n)=\widetilde{g}_{N}(n)+O(d(n)^{C}1_{\exists\,p\in\mathcal{J}_{N}\colon\,\,p\mid n}).$
Hence, by Shiu’s bound, for any $1\leq Y_{1}<Y_{2}\leq N$ with $Y_{2}\geq Y_{1}+N^{1/2}$ we can estimate
(5.13) $\mathbb{E}_{Y_{1}<n\leq Y_{2}}|\widetilde{g}(n)-\widetilde{g}_{N}(n)|\ll\sum_{p\in\mathcal{J}_{N}}\frac{1}{p}\mathbb{E}_{Y_{1}/p<m\leq Y_{2}/p}d(pm)^{C}\ll_{A}(\log N)^{2^{C}-1-A}.$
Hence, it suffices to prove (5.12) with $\widetilde{g}_{N}$ in place of $\widetilde{g}$.
For $I$ an interval, let $\omega_{I}(n)$ denote the number of prime factors from $I$ without multiplicities. Then we immediately have the Ramaré identity
$\widetilde{g}_{N}(n)=\sum_{p\in(Q_{N},R_{N})}\,\,\sum_{n=pm}\frac{g(pm)}{\omega_{(Q_{N},R_{N})}(pm)}.$
By multiplicativity, $g(pm)=g(p)g(m)$ and $\omega_{(Q_{N},R_{N})}(pm)=1+\omega_{(Q_{N},R_{N})}(m)$ unless $p\mid m$, so
$\widetilde{g}_{N}(n)$ $=\sum_{p\in(Q_{N},R_{N})}\,\,\sum_{n=pm}\frac{g(p)g(m)}{\omega_{(Q_{N},R_{N})}(m)+1}+O(1_{\exists\,p\in(Q_{N},R_{N})\colon\,\,p^{2}\mid n})$
$\coloneqq g^{\prime}_{N}(n)+O(1_{\exists\,p\in(Q_{N},R_{N})\colon\,\,p^{2}\mid n}).$
We trivially have
(5.14) $\mathbb{E}_{N-N(\log N)^{-A}<n\leq N}1_{\exists\,p\in(Q_{N},R_{N})\colon\,\,p^{2}\mid n}\ll\sum_{p\in(Q_{N},R_{N})}\frac{1}{p^{2}}\ll\frac{1}{Q_{N}}.$
Since $g^{\prime}_{N}(n)$ is by definition of the form $\sum_{n=ab}\alpha_{a}\beta_{b}$ with $|\alpha_{a}|\leq 1$, $|\beta_{b}|\ll d(b)^{C}$, and since $\alpha_{a},\beta_{a}$ are supported on $(Q_{N},N/Q_{N})$ by Lemma 5.8 and (5.14) we conclude that
$\left|\mathbb{E}_{N-N(\log N)^{-A}<n\leq N}\widetilde{g}_{N}(n)e(P(n))\right|\leq(\log N)^{-A}$
unless for some integer $1\leq\ell\ll_{A}(\log N)^{O_{A}(1)}$ the coefficients of $P$ satisfy
(5.15) $\|\ell c_{j}\|\ll_{A,C,k}\frac{(\log N)^{O_{A,C,k}(1)}}{N^{j}}$
for all integers $11\leq j\leq s-$ 1.
Suppose that (5.15) holds. Let $B$ be large in terms of $A,C,k$. Then $e(P(n))=e(P(N))(1+O((\log N)^{-A}))$ for $n$ belonging to any subinterval of $[1,N]$ of length $N(\log N)^{-A}$. Now, splitting into progressions modulo $\ell$ and into short intervals, we see that
$\left|\mathbb{E}_{N-N(\log N)^{-A}<n\leq N}\widetilde{g}_{N}(n)e(P(n))\right|$
$\ll$ $\ell\max_{b\ (\mathrm{mod}\ \ell)}\,\,\max_{z\in(N-N(\log N)^{-A},N)}\left|\mathbb{E}_{z-N(\log N)^{-B}<n\leq z}\widetilde{g}_{N}(n)e(P(n))1_{n\equiv b\ (\mathrm{mod}\ \ell)}\right|$
$\ll$ $\ell\max_{b\ (\mathrm{mod}\ \ell)}\,\,\max_{z\in(N-N(\log N)^{-A},N)}\left|\mathbb{E}_{z-N(\log N)^{-B}<n\leq z}\widetilde{g}_{N}(n)1_{n\equiv b\ (\mathrm{mod}\ \ell)}\right|+O_{A}((\log N)^{-A}).$
Now the claim follows from (5.13) and the Siegel–Walfisz assumption on $\widetilde{g}$ (which we have thanks to Lemma 5.7). ∎
6. Lemmas for the main proofs
@@ 6.1. A lacunary subsequence trick
We use a lacunary subsequence trick in the proofs of our pointwise convergence results. Such a trick roughly states that if $A_{N}\colon X\to\mathbb{C}$ are some measurable functions and we have strong quantitative decay for $\|A_{N_{j}}\|_{L^{1}(X)}$ for some lacunary sequence $(N_{j})_{j\in\mathbb{N}}$, then provided that $A_{N}$ does not vary too much on intervals of the form $[N_{j},N_{j+1}]$ in $L^{1}(X)$ norm, the sequence $A_{N}$ must converge to 00 in $L^{\infty}(X)$ norm. Variants of this idea are frequently used to establish convergence of ergodic averages; see for example [12, Section 5].
@@
Let $B\geq 40$. Let $(X,\nu)$ be a probability space, and for $N\geq 1$ let $A_{N}\colon X\to\mathbb{C}$ be a measurable function. If for any $N\geq 3$ we have
@@
and for $3\leq N<M\leq N(1+(\log N)^{-(B/2-1)})$ and for $\nu$-almost all $x\in X$ we have
(6.2) $|A_{M}(x)-A_{N}(x)|\leq(\log N)^{B/10-B/2+1},$
then for $\nu$-almost all $x\in X$ we have $\lim_{N\to\infty}(\log N)^{2B/5-2}|A_{N}(x)|=0$.
@@
Let $\eta>0$ be a small enough constant, and set $N_{j}=\exp(\eta\cdot j^{2/B})$ for all $j\in\mathbb{N}$. Note that $N_{j+1}\leq N_{j}(1+(\log N_{j})^{-(B/2-1)})$ for all large enough $j\in\mathbb{N}$, so by (6.2) we have
(6.3) $\sup_{M\in[N_{j},N_{j+1}]}|A_{M}(x)-A_{N}(x)|\leq(\log N_{j})^{B/10-B/2+1}$
For $\delta\in(0,1/2)$, denote
(6.4) $\begin{split}E(\delta)&=\{x\in X\colon\limsup_{N\to\infty}\,(\log N)^{2B/5-2}|A_{N}(x)|\geq\delta\},\\ E_{j}(\delta)&=\{x\in X\colon|A_{N_{j}}(x)|\geq\frac{\delta}{2}(\log N_{j})^{-2B/5+2}\}.\end{split}$
Then by (6.3) we have
(6.5) $E(\delta)\subset\bigcap_{i=3}^{\infty}\bigcup_{j\geq i}E_{j}(\delta).$
By Markov’s inequality and (6.1), we have the bound
$\nu(E_{j}(\delta))$ $\ll_{\delta}(\log N_{j})^{2B/5-B-2}\ll j^{-1.1}.$
Since $\sum_{j\geq 3}j^{-1.1}<\infty$, from (6.5) and the Borel–Cantelli lemma we conclude that $\nu(E(\delta))=0$. By the countable additivity of $\nu$, this then implies that
$\nu\left(\bigcap_{\delta>0}E(\delta)\right)=\nu\left(\bigcap_{j\geq 1}E(1/j)\right)=0,$
as desired. ∎